# A Multi-Agent Rollout Approach for Highway Bottleneck Decongenston in Mixed Autonomy

### v1

@@ Abstract
The integration of autonomous vehicles (AVs) into the existing transportation infrastructure offers a promising solution to alleviate congestion and enhance mobility. This research explores a novel approach to traffic optimization by employing a multi-agent rollout approach within a mixed autonomy environment. The study concentrates on coordinating the speed of human-driven vehicles by longitudinally controlling AVs, aiming to dynamically optimize traffic flow and alleviate congestion at highway bottlenecks in real-time. We model the problem as a decentralized partially observable Markov decision process (Dec-POMDP) and propose an improved multi-agent rollout algorithm. By employing agent-by-agent policy iterations, our approach implicitly considers cooperation among multiple agents and seamlessly adapts to complex scenarios where the number of agents dynamically varies. Validated in a real-world network with varying AV penetration rate and traffic flow, the simulations demonstrate that the multi-agent rollout algorithm significantly enhances performance, reducing average travel time on bottleneck segments by 9.42% with a 10% AV penetration rate.
I INTRODUCTION
@@
To address this gap, our study formulates the control tasks of multiple AVs as a decentralized partially observable Markov decision process (Dec-POMDP) and presents a novel multi-agent rollout approach. This approach reframes the multi-agent decision-making problem into a sequential decision-making paradigm using the agent-by-agent policy iteration (A2PI), and updates each AV’s policy via a Proximal Policy Optimization (PPO) algorithm [12]. By enabling the actions of preceding AVs to inform the state inputs of subsequent AVs, this method fosters an inherently adaptive cooperative interaction model. Crucially, our approach is designed to flexibly accommodate dynamic fluctuations in the number of AVs, ensuring robust performance in evolving traffic conditions. The main contributions of this article can be summarized as follows:
@@
The remainder of this paper is organized in the following manner. Section II introduces the model for the scenario depicted in Fig. 1. Section III details the enhanced multi-agent rollout approach. Section IV analyzes the performance of the proposed approach through simulations based on real traffic scenarios. Finally, Section V summarizes the main findings and outlines several directions for future work.
II Modeling and Formulation
@@
III Multi-agent Rollout Approach for Traffic Flow Optimization
@@
III-A Agent-by-agent policy iteration
@@
$J_{\bm{\pi}^{k}}(\bm{s}^{t})$ $=\mathbb{E}\left[r(\bm{s}^{t},a^{k,t}_{1},a^{k,t}_{2})+\gamma\phi_{(\pi^{k}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$\leq\max_{a_{1}^{t}\in\mathcal{A}_{1}}\mathbb{E}\left[r(\bm{s}^{t},a_{1}^{t},a^{k,t}_{2})+\gamma\phi_{(\pi^{k}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$=\mathbb{E}\left[r(\bm{s}^{t},a_{1}^{k+1,t},a^{k,t}_{2})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$\leq\max_{a_{2}^{k}\in\mathcal{A}_{2}}\mathbb{E}\left[r(\bm{s}^{t},a_{1}^{k+1,t},a_{2}^{k})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$=\mathbb{E}\left[r(\bm{s}^{t},a_{1}^{k+1,t},a_{2}^{k+1,t})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k+1}_{2})}(\bm{s}^{t+1})\right]$
@@
III-B Multi-agent rollout algorithm
@@
$L(\pi_{i}^{k})=\hat{\mathbb{E}}_{t}\Big{[}\min\Big{(}r_{t}(\pi_{i}^{k})\hat{A}_{t},\text{clip}\big{(}r_{t}(\pi_{i}^{k}),1-\epsilon,1+\epsilon\big{)}\hat{A}_{t}\Big{)}\Big{]},$ (6)
@@
IV Numerical Results
IV-A Experiment setting
@@ IV-B Compared methods
IV-C Performance comparison
Table I presents the performance of various methods across two traffic scenarios. In the first scenario, characterized by congestion with a traffic flow of 2400 $veh/hour$ including 6 AVs, the multi-agent rollout algorithm reduces the average travel time by 45.13% after two iterations compared to scenarios without control. Additionally, the final total reward obtained by the multi-agent rollout algorithm is higher than those achieved by MATD3 and MAPPO. In the second scenario, where traffic flow varies randomly from 0 to 4000 $veh/hour$, including a 10% AV penetration, the algorithm demonstrates a 9.42% improvement in average travel time compared to uncontrolled conditions. Here too, the final reward of our method surpasses those of MATD3 and MAPPO.
TABLE I: Performance comparisons of different algorithms
@@ IV-C Performance comparison
a T denotes simulation time; Perf denotes performance improvement;
   MARollout denotes multi-agent rollout algorithm
@@
Fig. 4 illustrates the reward curves of different algorithms in the first traffic scenario. It is observed that after two iterations, the MARollout achieves the highest reward. The evolution of rewards over the two iterations of the multi-agent rollout algorithm indicates the development of cooperation among agents. Initially, the total reward for all agents begins at a relatively low baseline within a stochastic environment. By the second iteration, there is a significant increase in the initial reward, exceeding the final reward of the first iteration. This improvement suggests that the agents developed effective policies during the first iteration to coordinate their actions. Fig. 5 illustrates this cooperative. We selected four vehicles near the bottleneck at a movement and depicted their action variations in trajectory form over the first six time steps. Among the three AVs, AV $2$ was passing through the bottleneck unaffected by other AVs, with its acceleration gradually increasing. Through observation, AV $1$ identified only one vehicle ahead in its lane and has no competition with AV $2$, so executing acceleration actions. Due to AV $3$ observing the presence of AV $2$ at the bottleneck, it decelerated in advance to avoid waiting.

### v2

# A Multi-Agent Rollout Approach for Highway Bottleneck Decongestion in Mixed Autonomy

@@ Abstract
The integration of autonomous vehicles (AVs) into the existing transportation infrastructure offers a promising solution to alleviate congestion and enhance mobility. This research explores a novel approach to traffic optimization by employing a multi-agent rollout approach within a mixed autonomy environment. The study concentrates on coordinating the speed of human-driven vehicles by longitudinally controlling AVs, aiming to dynamically optimize traffic flow and alleviate congestion at highway bottlenecks in real-time. We model the problem as a decentralized partially observable Markov decision process (Dec-POMDP) and propose an improved multi-agent rollout algorithm. By employing agent-by-agent policy iterations, our approach implicitly considers cooperation among multiple agents and seamlessly adapts to complex scenarios where the number of agents dynamically varies. Validated in a real-world network with varying AV penetration rates and traffic flow, the simulations demonstrate that the multi-agent rollout algorithm significantly enhances performance, reducing average travel time on bottleneck segments by 9.42% with a 10% AV penetration rate.
Keywords: Autonomous vehicles, traffic flow control, multi-agent rollout, reinforcement learning.
1 INTRODUCTION
@@
To address this gap, our study formulates the control tasks of multiple AVs as a decentralized partially observable Markov decision process (Dec-POMDP) and presents a novel multi-agent rollout approach. This approach reframes the multi-agent decision-making problem into a sequential decision-making paradigm using the agent-by-agent policy iteration (A2PI) [2], and updates each AV’s policy via a Proximal Policy Optimization (PPO) algorithm [10]. By enabling the actions of preceding AVs to inform the state inputs of subsequent AVs, this method fosters an inherently adaptive cooperative interaction model. Crucially, our approach is designed to flexibly accommodate dynamic fluctuations in the number of AVs, ensuring robust performance in evolving traffic conditions. The main contributions of this article can be summarized as follows:
@@
The remainder of this paper is organized in the following manner. Section 2 introduces the model for the scenario depicted in Fig. 1. Section 3 details the enhanced multi-agent rollout approach. Section 4 analyzes the performance of the proposed approach through simulations based on real traffic scenarios. Finally, Section 5 summarizes the main findings and outlines several directions for future work.
2 Modeling and Formulation
@@
3 Multi-agent Rollout Approach for Traffic Flow Optimization
@@
3.1 Agent-by-agent policy iteration
@@
$J_{\bm{\pi}^{k}}(\bm{s}^{t})$ $=\mathbb{E}\left[r^{t}(\bm{s}^{t},a^{k,t}_{1},a^{k,t}_{2})+\gamma\phi_{(\pi^{k}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$\leq\max_{a_{1}^{t}\in\mathcal{A}_{1}}\mathbb{E}\left[r^{t}(\bm{s}^{t},a_{1}^{t},a^{k,t}_{2})+\gamma\phi_{(\pi^{k}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$=\mathbb{E}\left[r^{t}(\bm{s}^{t},a_{1}^{k+1,t},a^{k,t}_{2})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$\leq\max_{a_{2}^{t}\in\mathcal{A}_{2}}\mathbb{E}\left[r^{t}(\bm{s}^{t},a_{1}^{k+1,t},a_{2}^{t})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k}_{2})}(\bm{s}^{t+1})\right]$
$=\mathbb{E}\left[r^{t}(\bm{s}^{t},a_{1}^{k+1,t},a_{2}^{k+1,t})+\gamma\phi_{(\pi^{k+1}_{1},\pi^{k+1}_{2})}(\bm{s}^{t+1})\right]$
@@
3.2 Multi-agent rollout algorithm
@@
$L(\pi_{i}^{k})=\hat{\mathbb{E}}_{t}\Big{[}\min\Big{(}r_{t}(\pi_{i}^{k})\hat{A}_{t},\text{clip}\big{(}r_{t}(\pi_{i}^{k}),1-\epsilon,1+\epsilon\big{)}\hat{A}_{t}\Big{)}\Big{]},$
@@
4 Numerical Results
4.1 Experiment setting
@@ 4.2 Compared methods
4.3 Performance comparison
Table 1 presents the performance of various methods across two traffic scenarios. In the first scenario, characterized by congestion with a traffic flow of 2400 $veh/hour$ including 6 AVs, the multi-agent rollout algorithm reduces the average travel time by 45.13% after two iterations compared to scenarios without control. Additionally, the final total reward obtained by the multi-agent rollout algorithm is higher than those achieved by MATD3 and MAPPO. In the second scenario, where traffic flow varies randomly from 0 to 4000 $veh/hour$, including a 10% AV penetration, the algorithm demonstrates a 9.42% improvement in average travel time compared to uncontrolled conditions. Here too, the final reward of our method surpasses those of MATD3 and MAPPO.
Table 1: Performance comparisons of different algorithms
@@ 4.3 Performance comparison
$\quad\quad\quad\quad\quad\quad{}^{a}$ T denotes simulation time;
$\quad\quad\quad\quad\quad\quad{}^{b}$erf denotes performance improvement;
$\quad\quad\quad\quad\quad\quad{}^{c}$ MARollout denotes multi-agent rollout algorithm
Fig.  4 illustrates the reward curves of different algorithms in the first traffic scenario. It is observed that after two iterations, the MARollout achieves the highest reward. The evolution of rewards over the two iterations of the multi-agent rollout algorithm indicates the development of cooperation among agents. Initially, the total reward for all agents begins at a relatively low baseline within a stochastic environment. By the second iteration, there is a significant increase in the initial reward, exceeding the final reward of the first iteration. This improvement suggests that the agents developed effective policies during the first iteration to coordinate their actions. Fig. 5 illustrates this cooperative. We selected four vehicles near the bottleneck at a movement and depicted their action variations in trajectory form over the first six time steps. Among the three AVs, AV $2$ was passing through the bottleneck unaffected by other AVs, with its acceleration gradually increasing. Through observation, AV $1$ identified only one vehicle ahead in its lane and has no competition with AV $2$, so executing acceleration actions. Due to AV $3$ observing the presence of AV $2$ at the bottleneck, it decelerated in advance to avoid waiting.